0 Required information The following information applies to the questions displayed below In 2018, the Westgate Construction Company entered into a contract to construct a road for Santa Clara County for $10,000,000. The road was completed in 2020. Information related to the contract is as follows: 2020 2019 $2,542,000 $3,772,000 $2,074,600 1,886, 000 2,020,000 4,294,000 3,686,000 1,810,000 3,800,000 4,390,000 2018 Cost incurred during the year Estimated costs to complete as of year-end Billings during the year Cash collections during the year 5,658,000 Westgate recognizes revenue over time according to percentage of completion.
Westgate recognizes revenue over time according to percentage of completion 4. Calculate the amount of revenue and gross profit (loss) to be recognized in each of the three years assuming the following costs incurred and costs to complete information. (Do not round intermediate calculations and round your final answers to the nearest whole dollar amount. Loss amounts should be indicated with a minus sign.) 2019 2,542,000 $3,810,000$3,210,000 2018 2020 Cost incurred during the year Estimated costs to eomplete as of year-end s, 658,000 3, 110,000 Revenue Gross profit (loss)
Answer #1
2018
Cost incurred during the year A 2,542,000
Estimated Cost B 5,658,000
Total Estimated Cost
C=A+B
8,200,000
Contract Value D 10,000,000
Revenue for 2018
E=(A/C)*D
3,100,000
Gross Profit E-A 558,000
2019
Earlier Cost incurred A 2,542,000
Cost incurred during the year B 3,772,000
Estimated Cost C 1,886,000
Total Estimated Cost D=A+B+C 8,200,000
Contract Value D 10,000,000
Total Revenue Shoul be booked E=((A+B)/C)*D 7,700,000
Less: Revenue allready booked in 2018
F 3,100,000
Revenue for 2019 G=E-F 4,600,000
Gross Profit
G-B
828,000
2020
Contract Value A 10,000,000
Revenue Booked in 2018 B 3,100,000
Revenue Booked in 2019 C 4,600,000
Revenue for 2020 D=A-B-C 2,300,000
Cost Incurred E 2,074,600
Gross Profit D-E 225,400
2018 2019 2020
Revenue $3,100,000 $4,600,000 $2,074,600
Gross profit $558,000 $828,000 $225,400